Battery acid is 4.22M aqueous H2SO4 soln and has density of 1.21 g/cm. What is the molarity of H2SO4 

Grade:12

1 Answers

Arun
5 years ago
This is a 4.22M solution.
1000cm³  of the solution contains 4.22 moles of H2SO4
Molar mass of H2SO4 = 1x2 + 32 + 16x4 = 98
Mass of H2SO4 in the solution = 4.22x98 = 413.56g
Total mass of the solution = 1000x1.21 = 1210g
Mass of water = 1210 - 413.56 = 796.44 g = 0.79644kg
Molality = Moles of solute/ kg of solvent
∴ Molality = 4.22/0.79644 = 5.3mol/Kg